two angles are supplementary. one angle is 280 degres less than three times the other. find measures of the angles . supplementary angles add to 180 degrees
Answer by mananth(12270)

let one angle be x
the other is 3x-280.
they are supplementary. sum = 180 deg.
x+3x-280=180
4x-280=180
4x=280+180
4x=460
/4
x=115 deg. one angle
180-115= 65 deg second angle